Hello there! I am having trouble with rate problems and how to solve them. here is one of my problems:
Two computers working together can finish a search in 30 seconds. One of these computers can finish in 50 seconds. How long would it take the second computer to finish the same search if it were working alone?

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Two computers working together can finish a search in 30 seconds. One of these computers can finish in 50 seconds. How long would it take the second computer to finish the same search if it were working alone?
------
Together DATA::
time= 30 seconds/job ; rate = 1/30 job/second
---------------------
One computer DATA:
time = 50 seconds/job ; rate = 1/50 job/second
----
2nd computere DATA::
time = x seconds/job ; rate = 1/x job/second
-------------
Equation:
rate + rate = together rate
1/50 + 1/x = 1/30
-------
30x + 50*30 = 50x
----
20x = 1500
x = 75 minutes (time of 2nd computer to do the job)
-------------
Cheers,
Stan H.
--------

Answer by ikleyn(53765)

.
The correct answer is 75 seconds.